OpenStudy (lolipop_drum312):

What is the slope of the line passing through the points (2, −5) and (4, 1)? A. 3 B.-4/5 C. 2 D. 5/4

563blackghost (563blackghost):

Use the slope formula... \(\huge{\frac{y_{2}-y_{1}}{x_{2}-x_{1}} \rightarrow {\frac{1-(-5)}{4-2}}}\)

OpenStudy (narutofan2016):

a

OpenStudy (lolipop_drum312):

ok are you sure tho

OpenStudy (quickstudent):

To find the slope, use the slope formula: m = (y2 - y1)/(x2 - x1) m = (1 - -5)/(4 - 2) m = 6/2 m = 3 You can also use a quick way with this calculator: http://calculator.tutorvista.com/slope-calculator.html

OpenStudy (narutofan2016):

ya i am 1- -5 would be 1+5 then you get 6 so 4-2 would be 2 6 divided by 2 would equal 3

OpenStudy (lolipop_drum312):

ok thank you both so who wants the medal or like can one of ya'll medal one another; i will emdal one of you guys

OpenStudy (narutofan2016):

whichever one you want to give it to

OpenStudy (lolipop_drum312):

ok i gave it to you NF2016 can you give the other one, one too

OpenStudy (narutofan2016):

okay

OpenStudy (lolipop_drum312):

ok here's another one

OpenStudy (lolipop_drum312):

should i open up new panle or just post it here

OpenStudy (quickstudent):

YW and Thank you for the medal:) And open the question in new post:)

OpenStudy (lolipop_drum312):

ok i will and you are welcome my friend *peace sign*